OpenStudy (anonymous):

Not understanding this concept, please Help! Find the inverse formula for the following function: See the first reply for the equation I need to find an inverse for.

OpenStudy (anonymous):

\(\ \huge y = \frac{1+e^x}{1-e^x}. Find the inverse? \)

OpenStudy (anonymous):

The attempt at a solution \[(1+e^x/1-e^x) (1+e^x/1+e^x) = (2+e^x^2/2)\] \[\ln(2+e^x^2/2)= \ln2+x^2/\ln2\]

OpenStudy (anonymous):

The attempt at a solution 1) (1+e^x/1-e^x)(1+e^x/1+e^x)=(2+e^x^2/2) ln(2+e^x^2/2)= ln2+x^2/ln2

OpenStudy (anonymous):

@Study23 is that correct!

OpenStudy (anonymous):

So did you switch the y and the x?

OpenStudy (anonymous):

you solve for x then at the end you swap variables.

OpenStudy (anonymous):

So how do you begin to solve for x?

OpenStudy (anonymous):

if the question was this:\[y=\frac{1+z}{1-z}\]would you be able to solve for z?

OpenStudy (anonymous):

because if you can do this, your problem requires one more step afterwards.

OpenStudy (anonymous):

x = ln((-1+y)/(1+y)) now you can swap x for y

OpenStudy (anonymous):

Well wouldn't you multiply both sides by (1-z), @joemath314159 ?

OpenStudy (anonymous):

@timo86m how did you get to that x=ln part?

OpenStudy (anonymous):

yes, thats correct:\[y=\frac{1+z}{1-z}\iff y(1-z)=1+z\iff y-yz=1+z\]\[\iff -z-yz=1-y\iff z(-1-y)=1-y\iff \]\[z=\frac{1-y}{-1-y}\]

OpenStudy (anonymous):

Now switch z with e^x, and it should be clear how to proceed.

OpenStudy (anonymous):

So, @joemath314159 basically you substituted a variable \(\ \huge \color{red}{z}\) for \(\ \huge \color{green}{e^x}?\)

OpenStudy (anonymous):

Wouldn't you have to use \(\ \huge \color{blue}{ln}?\).

OpenStudy (anonymous):

yes. its not necessary, but for some reason the problem looks sorta hard with e^x's floating about. yes you would have to use ln after where I left off.

OpenStudy (anonymous):

Where would use ln? Where would you substitute the z value

OpenStudy (anonymous):

@joemath314159

OpenStudy (anonymous):

apply ln of both sides of jomaths final equation solved for z

OpenStudy (anonymous):

since z= e^x the ln of e^x is ln^(e^x) is just x so x= ln ((1-y)/(-1-y))